# Improving exponential-family random graph models for bipartite networks

### v1

@@ Abstract
Bipartite graphs, representing two-mode networks, arise in many research fields. These networks have two disjoint node sets representing distinct entity types, for example persons and groups, with edges representing associations between the two entity types. In bipartite graphs, the smallest possible cycle is a cycle of length four, and hence four-cycles are the smallest structure to model closure in such networks. Exponential-family random graph models (ERGMs) are a widely used model for social, and other, networks, including specifically bipartite networks. Existing ERGM terms to model four-cycles in bipartite networks, however, are relatively rarely used. In this work we demonstrate some problems with these existing terms to model four-cycles, and define new ERGM terms to help overcome these problems. The position of the new terms in the ERGM dependence hierarchy, and their interpretation, is discussed. The new terms are demonstrated in simulation experiments, and their application illustrated with ERGM models of empirical networks ranging in size from hundreds of nodes to hundreds of thousands of nodes.
@@ 1 Introduction
Two-mode networks can be studied by means of their projections onto one-mode (unipartite) networks, thereby allowing the use of existing methods for one-mode networks, however this can result in lost information, and properties of the one-mode networks (such as high clustering coefficients) that are due to the projection process rather than the original data [4]. Although the former problem can be overcome by using both projections in analyses [5, 6], it is still desirable to directly study the original two-mode network, for which specific methods are required [4].
@@
In studying one-mode networks, a central concept is triadic closure, the tendency for a path of length two (a “two-path”; three nodes connected by two edges) to be “closed” into a triangle by the addition of a third edge. In the context of a social network, this is the process of a friend of a friend becoming themselves a friend, and is perhaps most well known via the “strength of weak ties” [7] argument, whereby an open two-path of strong ties is the “forbidden triad”, which is “forbidden” because the two actors with strong ties to a common third actor must themselves have a strong tie. In a bipartite network, however, a closed triad (triangle) is impossible; indeed it is a defining feature of bipartite graphs that only cycles of even length are possible [8]. Therefore, the smallest possible cycle in a bipartite graph is a four-cycle, and hence four-cycles are frequently used to measure closure in bipartite networks [9, 10].
@@
Exponential random graph models (ERGMs) are a widely used model for social [11, 12, 13, 14], and other [15, 16, 17] networks. Specific forms of the ERGM have been developed for two-mode networks [18, 19, 20, 21], however, as we shall show in this work, existing ERGM models for bipartite networks often have problems modelling four-cycles, and hence can frequently not adequately model closure in bipartite networks.
@@
In this work, we will show that, despite the importance of four-cycles in two-mode networks, ERGM terms to model four-cycles in such networks are relatively rarely used, in contrast to the ubiquity of terms modelling triadic closure in one-mode networks. We will then describe some problems with existing configurations for modelling four-cycles that could explain their relatively infrequent use, and propose new ERGM configurations for modelling four-cycles to help overcome these problems. We will discuss the interpretation of these new parameters, and their position in the dependence hierarchy of Pattison & Snijders [22]. We will then demonstrate the new configurations using simulation experiments and apply them in ERGM modelling of seven empirical networks in a variety of domains (ecology, criminology, and sociology) and ranging in size from hundreds of nodes to hundreds of thousands of nodes.
@@ 2 The exponential-family random graph model (ERGM)
ERGM models with only simple configurations (such as the Markov edge plus triangle model) can be prone to problems with phase transitions or “near-degeneracy” [29, 30, 31, 32, 33, 34, 35, 14]. Such problems are typically avoided by the use of “alternating” [30, 36, 37] or “geometrically weighted” [38, 39] configurations. The former are parameterised with a decay parameter $\lambda$ controlling the rate at which the weight of contributions from additional terms in the statistic decay. The corresponding parameter for the geometrically weighed configurations can be estimated as part of the model, in which case it becomes a “curved ERGM” [40], however in this work we will use fixed values of $\lambda$ for the “alternating” configurations.
@@ 3 Literature survey
In order to get an overview of which effects are used in modelling bipartite networks with ERGMs, and how well they fit four-cycles, we conducted a comprehensive survey of publications which included ERGM models of bipartite networks. To be included, a publication must contain one or more ERGM models of one or more empirical bipartite (two-mode) networks. There must be sufficient detail given to know at least the parameters included in the models, and their estimated signs and statistical significance. Models of one-mode projections of two-mode networks were excluded; we only consider ERGM models of the bipartite network itself. Models of multilevel networks were excluded, although if there is a model of just the cross-level (bipartite) network, this is included.
We used Google Scholar (search date 24 November 2022) to search for papers with terms “bipartite“ and “ergm”,“bpnet” and “ergm”, “statnet” and “bipartite”, and “gwb1dsp” or “gwb2dsp” (the latter being statnet [41, 42] terms used for bipartite networks). We also included papers that we had prior knowledge of being about bipartite ERGMs, and followed backward and forward citations (using Google Scholar for the latter) to find other relevant papers. The papers included in the survey are listed in Table LABEL:tab:literature (Appendix A).
@@
Table 1 summarises the parameter estimates in models from the literature in Table LABEL:tab:literature that were estimated using BPNet or MPNet, and contain the four-cycles parameter C4, or the bipartite alternating $k$-two-path parameters ($\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ in BPNet, XACA and XACB in MPNET) defined in Wang et al. [18]. Less than a third ($20/63$) include the four-cycles parameter, less than half ($30/63$) include either of the two alternating $k$-two-path parameters, and less than 30% ($17/63$) include both $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$.
@@
Table 2 summarises the parameter estimates in models from the literature in Table LABEL:tab:literature that were estimated using statnet, and which contain the four-cycle term, the bipartite geometrically weighted dyadwise shared partner distribution (gwb1dsp or gwb2dsp) term, the statnet equivalent of the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ parameters, or the geometrically weighted non-edgewise shared partner (gwnsp) term. The counts for gwb1dsp and gwb2dsp include those for the bipartite geometrically weighted non-edgewise shared partner terms gwb1nsp and gwb2nsp, which are equivalent to gwb1dsp and gwb2dsp, respectively, for bipartite networks. Only one model estimated with statnet contained an explicit term for four-cycles [55]. Further, only $5/43$ models contain a gwb1dsp pr gwb2dsp term at all, and only two models contain both (Lubell & Robbins [56] have two models, both of which include both gwb1nsp and gwb2nsp).
@@
This relative rarity of models containing terms to model closure (four-cycles) in bipartite networks, or assess goodness-of-fit to four-cycles, is in stark contrast to ERGM modelling for one-mode networks, where terms modelling triadic closure, such as triangles, alternating-$k$-triangles, or geometrically weighted edgewise shared partners (gwesp in statnet) are almost always included in models, since triadic closure (as evidenced by clustering, or transitivity, in the network), is a well-known feature of social networks [7, 57, 58]. For example, Clark & Handcock [59] use the latent order logistic (LOLOG) model [60] to reproduce ERGM models for 13 networks from peer reviewed papers, and all of these models contain alternating-$k$-triangles [36, 61, 62, 63], gwesp [64, 65, 66, 67, 68, 69], or three-cycles [61, 70] terms, and in the majority of cases the results are able to be replicated with a triangle term in the LOLOG model [59]. And yet, despite higher than expected numbers of four-cycles (or bipartite clustering coefficient) being a notable feature of some bipartite networks, such as director interlock networks [9, 10], and collaboration and communication networks [10], the literature survey presented here shows that terms to model four-cycles are relatively rarely used in published ERGM models of bipartite networks. This could be due to researchers choosing not to model bipartite closure, however, as shown by some examples in the following section, it can also be due to difficulties in obtaining converged model estimates when using existing model terms.
@@ 4 Problems with existing bipartite ERGM statistics
It is notable that in the original paper proposing $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ configurations [18], they are explicitly described as $k$-two-path statistics, and in fact they are just bipartite versions of the one-mode alternating-$k$-two-path statistic, representing multiple shared partners (gwdsp in statnet, and then later gwb1dsp and g2b2dsp for bipartite networks). There is, however, already some “semantic slippage” into interpreting them as “cycles” or “closure” — even though they only actually include cycles (closure) when $k>1$ ($k=1$ is a two-path, $k=2$ is a four-cycle; see Fig. 3 in Wang et al. [18]). For example, “… a better chance of achieving model convergence when closure effects ($K-C_{P}$ and $K-C_{A}$) are included in the model.” [18, p. 22]. Even the name $\mathrm{K\text{-}C_{A}}$ (or $\mathrm{K\text{-}C_{P}}$) suggests “cycle” or “closure” by the use of the “C” (C4 is used for four-cycles in the paper). Wang et al. [19] goes back to naming the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ statistics as A2P-A and A2P-B and describing them as “shared affiliations (alternating two-paths)” [19, p. 213]. However the MPNet terminology is XACA and XACB [44, 45], again with the “C” suggestive of cycles or closure, and the EstimNetDirected [71] software (https://github.com/stivalaa/EstimNetDirected) refers to these effects as BipartiteAltKCyclesA and BipartiteAltKCyclesB.
@@
In the influential book edited by Lusher et al. [11], Wang [20] describes $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ explicitly as “alternating A cycles” and “alternating P cycles” [20, p. 124], with [20, Fig. 10.11] captioned “Alternating 2-paths” but with the figure panels labelled “A cycles (KCA)” and “P cycles (KCP)” respectively [20, p. 125]. This description or interpretation is carried over into the empirical part of the book, with Harrigan & Bond [72], applying bipartite ERGM to a director interlock network, describing $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ as “alternating $k$-cycles” for directors and corporations respectively [72, p. 266], and in the ERGM results tables as “Director 4-cycles” and “Corporation 4-cycles” [72, pp. 268-269]. It is also notable that Harrigan & Bond [72] describes the difficulty of fitting models with $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ and the resulting poor goodness-of-fit for the four-cycle statistic:
@@
We found for this particular network that we cannot have both K-Cp and K-Ca in the same model due to convergence issues. In line with Wang, Sharpe, Robins, and Pattison (2009), we present two alternative models, one with each possible k-cycle parameter. … However, in line with Wang, Sharpe, Robins, and Pattison (2009), we had a poor fit on the classic 4-cycle parameter [C4], which suggests that improving these structural effects is a substantial area of future research.
[72, p. 267]
@@
Given the results of the literature review described in Section 3 (which includes Harrigan & Bond [72]) it would seem, however, that no such research improving these structural effects has been published yet, and this work may be the first attempt to do so.
@@
We may conclude from this that the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics count too many things other than four-cycles to be usefully used and interpreted as bipartite closure in many cases. Specifically, they count two-paths as their first (highest weighted) term. This results in situations where long paths or cycles contribute to the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ statistics, despite having exactly zero four-cycles (see Table 3). Perhaps even more problematically, high-degree nodes, or stars, for example the “Nine-star” structure in Table 3, result in large values of $\mathrm{K\text{-}C_{A}}$ or $\mathrm{K\text{-}C_{P}}$ (XACA or XACB in MPNet terminology), depending on which node set the hub node is in, but also have exactly zero four-cycles. It seems clear that large networks are likely to contain many stars, and many paths (of length two or more; and note that a two-star is just a two-path), as well as large cycles [73, 74], and these will contribute to large values of the $\mathrm{K\text{-}C_{P}}$ and/or $\mathrm{K\text{-}C_{A}}$ statistics, but nothing to the number of four-cycles (the C4 statistic).
@@
In addition, interpretation of the alternating $k$-two-paths parameter is problematic: “In this article, we do not concentrate on alternating $k$-two-path parameters. For some data, we have found it important to include them in models but further work is needed to understand better their effect when included with other parameters” [36, p. 201]. Martin [75] described this as the authors “being somewhat mystified by this statistic” [75, p. 86], and that most ERGM modellers would not be able to describe a “clear behavioral-process analogue to the once-canonical alternating two paths statistic” [75, p. 86].
@@
Therefore, we propose a new statistic that counts four-cycles (and not two-paths), but which is less prone to problems with near-degeneracy than the simple four-cycles parameter. (A simpler new statistic, based on $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ but which does not count two-paths, which turns out to be more prone to near-degeneracy than $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$, is described in Appendix B).

### v3

@@ Abstract
Bipartite graphs, representing two-mode networks, arise in many research fields. These networks have two disjoint node sets representing distinct entity types, for example persons and groups, with edges representing associations between the two entity types. In bipartite graphs, the smallest possible cycle is a cycle of length four, and hence four-cycles are the smallest structure to model closure in such networks. Exponential-family random graph models (ERGMs) are a widely used model for social, and other, networks, including specifically bipartite networks. Existing ERGM terms to model four-cycles in bipartite networks, however, are relatively rarely used. In this work we demonstrate some problems with these existing terms to model four-cycles, and define new ERGM terms to help overcome these problems. The position of the new terms in the ERGM dependence hierarchy, and their interpretation, is discussed. The new terms are demonstrated in simulation experiments, and their application illustrated on a canonical example of an empirical two-mode network.
@@ 1 Introduction
Two-mode networks can be studied by means of their projections onto one-mode (unipartite) networks, thereby allowing the use of existing methods for one-mode networks, however this can result in lost information, and properties of the one-mode networks (such as high clustering coefficients) that are due to the projection process rather than the original data [4]. Although the former problem can be ameliorated by using both projections in analyses [5, 6], it is still desirable to study the original two-mode network directly, for which specific methods are required [4].
@@
In studying one-mode networks, a central concept is triadic closure, the tendency for a path of length two (a “two-path”; three nodes connected by two edges) to be “closed” into a triangle by the addition of a third edge. In the context of social networks, this is the process of a friend of a friend becoming themselves a friend, and is perhaps most well known via the “strength of weak ties” [7] argument, whereby an open two-path of strong ties is the “forbidden triad”, which is “forbidden” because the two actors with strong ties to a common third actor must themselves have a strong tie. In a bipartite network, however, a closed triad (triangle) is impossible; indeed it is a defining feature of bipartite graphs that only cycles of even length are possible [8]. Therefore, the smallest possible cycle in a bipartite graph is a four-cycle, and hence four-cycles are frequently used to measure closure in bipartite networks [9, 10].
An example of a social process that creates four-cycles is peer referral in director interlock networks [9, 11]. If a director on the boards of two companies recruits a director from one of them to also sit on the board of the other, then an open three-path is closed, forming a four-cycle.
Four-cycles in bipartite networks also have the particular importance that, together with the degree distribution, they explain the degree assortativity in the one-mode projected network [12, 13].
@@
Exponential random graph models (ERGMs) are a widely used model for social [14, 15, 16, 17], and other [18, 19, 20] networks. Specific forms of the ERGM have been developed for two-mode networks [21, 22, 23, 24], however, as we shall show in this work, existing ERG models for bipartite networks often have problems modelling four-cycles, and hence can frequently not adequately model closure in bipartite networks.
@@
In this work, we will show that, despite the importance of four-cycles in two-mode networks, ERGM terms to model four-cycles in such networks are relatively rarely used, in contrast to the ubiquity of terms modelling triadic closure in one-mode networks. We will then describe some problems with existing configurations for modelling four-cycles that could explain their relatively infrequent use, and propose new ERGM configurations for modelling four-cycles to help overcome these problems. We will discuss the interpretation of these new parameters, and their position in the dependence hierarchy of Pattison & Snijders [25]. We will then demonstrate the new configurations using simulation experiments and demonstrate their use on a canonical example of an empirical two-mode network.
@@ 2 The exponential-family random graph model (ERGM)
ERG models with only simple configurations (such as the Markov edge plus triangle model) can be prone to problems with phase transitions or “near-degeneracy” [32, 33, 34, 35, 36, 37, 38, 17]. Such problems are typically avoided by the use of “alternating” [33, 39, 40] or “geometrically weighted” [41, 42] configurations. The former are parameterised with a decay parameter $\lambda$ controlling the rate at which the weight of contributions from additional terms in the statistic decay. The corresponding parameter for the geometrically weighed configurations can be estimated as part of the model, in which case it becomes a “curved ERGM” [43], however in this work we will use fixed values of $\lambda$ for the “alternating” configurations.
In order to model two-mode networks, and in particular affiliation networks, Wang et al. [21] define the alternating $k$-two-path statistics $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$, implemented as XACA and XACB in MPNet [44, 45], and BipartiteAltKCyclesA and BipartiteAltKCyclesB in EstimNetDirected (https://github.com/stivalaa/EstimNetDirected):
$z_{\mathrm{K\text{-}C_{A}}(\lambda)}=z_{\mathrm{XACA}(\lambda)}=z_{\text{BipartiteAltKCyclesA}(\lambda)}=\lambda\sum_{i\in B}\sum_{\left\{l\in B\,:\,l<i\right\}}\left[1-\left(1-\frac{1}{\lambda}\right)^{L_{2}(i,l)}\right]$ (2)
where $\lambda>1$ is the decay parameter, and $L_{2}(i,l)$ is the number of two-paths connecting $i$ to $l$:
$L_{2}(i,l)=\sum_{h\neq i,l}x_{ih}x_{hl}.$ (3)
The corresponding change statistic is [21]:
$\delta_{\mathrm{K\text{-}C_{A}}(\lambda)}(i,j)=\delta_{\mathrm{XACA}(\lambda)}(i,j)=\delta_{\text{BipartiteAltKCyclesA}(\lambda)}(i,j)$ $=\sum_{l\in B}\left[x_{il}\left(1-\frac{1}{\lambda}\right)^{L_{2}(j,l)}\right]$ (4)
$=\sum_{l\in N(i)}\left(1-\frac{1}{\lambda}\right)^{L_{2}(j,l)}.$ (5)
where $N(i)$ denotes the neighbours of node $i$, that is, nodes $k\neq i$ such that $x_{ik}=1$, or, equivalently, $d(i,k)=1$, where $d(u,v)$ is the geodesic distance from $u$ to $v$. The statistic and change statistic for BipartiteAltKCyclesB ($\mathrm{K\text{-}C_{P}}$ or XACB) are defined similarly.
@@ 3 Literature survey
In order to get an overview of which effects are used in modelling bipartite networks with ERGMs, and how well they fit four-cycles, we conducted a comprehensive survey of publications which included ERG models of bipartite networks. To be included, a publication must contain one or more ERG models of one or more empirical bipartite (two-mode) networks. There must be sufficient detail given to know at least the parameters included in the models, and their estimated signs and statistical significance. Models of one-mode projections of two-mode networks were excluded; we only consider ERG models of the bipartite network itself. Models of multilevel networks were excluded, although if there is a model of just the cross-level (bipartite) network, this is included. The papers included in the survey are listed in Table LABEL:tab:literature in Appendix A, which also contains further details of the literature survey.
@@
Table 1 summarises the parameter estimates in models from the literature in Table LABEL:tab:literature that were estimated using BPNet or MPNet, and contain the four-cycles parameter C4, or the bipartite alternating $k$-two-path parameters ($\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$) defined in Wang et al. [21]. Less than a third ($20/63$) include the four-cycles parameter, less than half ($30/63$) include either of the two alternating $k$-two-path parameters, and less than 30% ($17/63$) include both $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$.
@@
Table 2 summarises the parameter estimates in models from the literature in Table LABEL:tab:literature that were estimated using statnet, and which contain the four-cycle term, the bipartite geometrically weighted dyadwise shared partner distribution (gwb1dsp or gwb2dsp) term, the statnet equivalent of the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ parameters, or the geometrically weighted non-edgewise shared partner (gwnsp) term. Only one model estimated with statnet contained an explicit term for four-cycles [58]. Further, only $5/43$ models contain a gwb1dsp pr gwb2dsp term at all, and only two models contain both (Lubell & Robbins [59] have two models, both of which include both gwb1nsp and gwb2nsp).
@@
This relative rarity of models containing terms to model closure (four-cycles) in bipartite networks, or assess goodness-of-fit to four-cycles, is in stark contrast to ERG modelling for one-mode networks, where terms modelling triadic closure, such as triangles, alternating-$k$-triangles, or geometrically weighted edgewise shared partners (gwesp in statnet) are almost always included in models, since triadic closure (as evidenced by clustering, or transitivity, in the network), is a well-known feature of social networks [7, 60, 61]. For example, Clark & Handcock [62] use the latent order logistic (LOLOG) model [63] to reproduce ERG models for 13 networks from peer reviewed papers, and all of these models contain alternating-$k$-triangles [39, 64, 65, 66], gwesp [67, 68, 69, 70, 71, 72], or three-cycles [64, 73] terms, and in the majority of cases the results are able to be replicated with a triangle term in the LOLOG model [62]. And yet, despite higher than expected numbers of four-cycles (or bipartite clustering coefficient) being a notable feature of some bipartite networks, such as director interlock networks [9, 10], and collaboration and communication networks [10], the literature survey presented here shows that terms to model four-cycles are relatively rarely used in published ERG models of bipartite networks. This could be due to researchers choosing not to model bipartite closure, however, as shown by some examples in the following section, it can also be due to difficulties in obtaining converged model estimates when using existing model terms.
@@ 4 Problems with existing bipartite ERGM statistics
It is notable that in the original paper proposing $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ configurations [21], they are explicitly described as $k$-two-path statistics, and in fact they are just bipartite versions of the one-mode alternating-$k$-two-path statistic, representing multiple shared partners (gwdsp in statnet, and then later gwb1dsp and gwb2dsp for bipartite networks). There is, however, already some “semantic slippage” into interpreting them as “cycles” or “closure” — even though they only actually include cycles (closure) when $k>1$ ($k=1$ is a two-path, $k=2$ is a four-cycle; see Fig. 3 in Wang et al. [21]). For example, “… a better chance of achieving model convergence when closure effects ($K-C_{P}$ and $K-C_{A}$) are included in the model.” [21, p. 22]. Even the name $\mathrm{K\text{-}C_{A}}$ (or $\mathrm{K\text{-}C_{P}}$) suggests “cycle” or “closure” by the use of the “C” (C4 is used for four-cycles in the paper). Wang et al. [22] goes back to naming the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ statistics as A2P-A and A2P-B and describing them as “shared affiliations (alternating two-paths)” [22, p. 213]. However the MPNet terminology is XACA and XACB [44, 45], again with the “C” suggestive of cycles or closure, and the EstimNetDirected [74] software (https://github.com/stivalaa/EstimNetDirected) refers to these effects as BipartiteAltKCyclesA and BipartiteAltKCyclesB.
@@
In the influential book edited by Lusher et al. [14], Wang [23] describes $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ explicitly as “alternating A cycles” and “alternating P cycles” [23, p. 124], with [23, Fig. 10.11] captioned “Alternating 2-paths” but with the figure panels labelled “A cycles (KCA)” and “P cycles (KCP)” respectively [23, p. 125]. This description or interpretation is carried over into the empirical part of the book, with Harrigan & Bond [75], applying bipartite ERGM to a director interlock network, describing $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ as “alternating $k$-cycles” for directors and corporations respectively [75, p. 266], and in the ERGM results tables as “Director 4-cycles” and “Corporation 4-cycles” [75, pp. 268-269]. It is also notable that Harrigan & Bond [75] describes the difficulty of fitting models with $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ and the resulting poor goodness-of-fit for the four-cycle statistic:
@@
We found for this particular network that we cannot have both K-Cp and K-Ca in the same model due to convergence issues. In line with Wang, Sharpe, Robins, and Pattison (2009), we present two alternative models, one with each possible k-cycle parameter. … However, in line with Wang, Sharpe, Robins, and Pattison (2009), we had a poor fit on the classic 4-cycle parameter [C4], which suggests that improving these structural effects is a substantial area of future research. [75, p. 267]
@@
Given the results of the literature review described in Section 3 it would seem, however, that no such research improving these structural effects has been published yet, and this work may be the first attempt to do so.
@@
We may conclude from this that the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics count too many things other than four-cycles to be usefully used and interpreted as bipartite closure in many cases. Specifically, they count simple two-paths ($k$-two-paths with $k=1$) as their first (highest weighted) term. This results in situations where long paths or cycles contribute to the $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ statistics, despite having exactly zero four-cycles (see Table 3). Perhaps even more problematically, high-degree nodes, or stars, for example the “Nine-star” structure in Table 3, result in large values of $\mathrm{K\text{-}C_{A}}$ or $\mathrm{K\text{-}C_{P}}$ (XACA or XACB in MPNet terminology), depending on which node set the hub node is in, but also have exactly zero four-cycles. It seems clear that large networks are likely to contain many stars, and many paths (of length two or more; and note that a two-star is just a two-path), as well as large cycles [76, 77], and these will contribute to large values of the $\mathrm{K\text{-}C_{P}}$ and/or $\mathrm{K\text{-}C_{A}}$ statistics, but nothing to the number of four-cycles (the C4 statistic).
@@
In addition, Robins et al. [39] report that interpretation of the alternating $k$-two-paths parameter is problematic: “In this article, we do not concentrate on alternating $k$-two-path parameters. For some data, we have found it important to include them in models but further work is needed to understand better their effect when included with other parameters” [39, p. 201]. Martin [78] described this as the authors “being somewhat mystified by this statistic” [78, p. 86], and that most ERG modellers would not be able to describe a “clear behavioral-process analogue to the once-canonical alternating two paths statistic” [78, p. 86].
@@
Therefore, we propose new statistics that count four-cycles (and not two-paths). In the following two sections we propose two different new statistics. In Section 5 we propose a simple new statistic based on $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$. Unfortunately, however, this statistic is shown to be more, rather than less, prone to degeneracy than $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$. Therefore, in Section 6 we propose a new statistic based on counting four-cycles, but which is less prone to problems with near-degeneracy than the simple four-cycles parameter or the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters.
5 A simple, but unsuccessful, new statistic
A simple solution to the problem, described in Section 4, that the $\mathrm{K\text{-}C_{A}}$ (2) and $\mathrm{K\text{-}C_{P}}$ statistics count simple two-paths as their first term ($k$-two-paths, $k=1$), is to remove the first term and reverse the signs. In this way, the first, positive, term no longer counts open two-paths, but rather counts four-cycles. We define a new statistic BipartiteAltK4CyclesA as
and its change statistic
where
is the number of two-paths connecting nodes in node set $B$ (and which therefore go through a node in node set $A$), $L_{2}(i,l)$, defined by (3), is the number of two-paths connecting $i$ to $l$, and $\operatorname{deg}(i)$, the degree of node $i$, is the change statistic for the number of two-paths through node $i$. BipartiteAltK4CyclesB and its change statistic are defined similarly.
Table 4 is a copy of Table 3, but with the new BipartieAltK4CyclesA and BipartiteAltK4CyclesB statistics (labelled BpAK4CA and BpAK4CB respectively) included (and the $\mathrm{N_{A}}$, $\mathrm{N_{B}}$, L, BpNP4CA, and BpNP4CB columns removed to make space). This table shows the value of the new statistics on some small example networks, demonstrating, that, by design, they are zero for networks in which there are no four-cycles.
Table 4: Statistics of some example bipartite networks. C4 is the number of four-cycles, and BpAK4CA and BpAK4CB are the new statistics BipartiteAltK4CyclesA and BipartiteAltK4CyclesB, respectively. Nodes in node set A are represented as red circles, and nodes in node set B as blue squares.
Name Visualization C4 XACA XACB BpAK4CA BpAK4CB
Two-path [Uncaptioned image] 0 1 0 0 0
Four-cycle [Uncaptioned image] 1 1.5 1.5 0.5 0.5
Four-cycles-3 [Uncaptioned image] 3 4.5 1.75 1.5 1.25
Ten-cycle [Uncaptioned image] 0 5 5 0 0
Nine-star [Uncaptioned image] 0 36 0 0 0
Four-fan-3 [Uncaptioned image] 3 16.5 4.5 1.5 1.5
Unfortunately, however, simulation experiments indicate that the new BipartiteAltK4CyclesA and BipartiteAltK4CyclesB parameters are actually more problematic with respect to near-degeneracy than the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters. Fig. 1 shows the results of simulation experiments similar to those described by Wang et al. [21, p. 19]. Bipartite networks were simulated with 30 nodes in node set $A$ and 20 nodes in node set $B$ with the Edge parameter set to $-3.0$. In three different sets of simulations, for each of the BipartiteAltKCyclesB, BipartiteAltK4CyclesB, and BipartiteFourCyclesNodePowerB (defined in Section 6) parameters, the parameter in question is varied from $-1.00$ to $10.0$ in increments of $0.01$ for each of two values of $\lambda$ ($\lambda=2$ and $\lambda=5$) for BipartiteAltKCyclesB and BipartiteAltK4CyclesB, and for each of two values of $\alpha$ ($\alpha=1/2$ and $\alpha=1/5$) for BipartiteFourCyclesNodePowerB. The networks were simulated using the SimulateERGM program from the EstimNetDirected software package, using the basic ERGM sampler, with a burn-in of $10^{5}$ iterations and an interval of $10^{4}$ iterations between each of 100 samples, to ensure that samples are drawn from the equilibrium ERGM distribution, and are not too autocorrelated.
Figure 1: Effect of varying (a) the BipartiteAltKCyclesB parameter, (b) the BipartiteAltK4CyclesB parameter, and (c) the BipartiteFourCyclesNodePowerB parameter on the statistic corresponding to the parameter itself (left) and the Edge statistic (right). Each graph plots the mean value of the statistic (over 100 simulations) for two different values of the relevant $\lambda$ or $\alpha$ parameter.
The right panel in Fig. 1(a) shows the results of a simulation similar to that shown in Fig, 9 of Wang et al. [21]: the number of edges increases smoothly, giving good coverage of the graph space (with respect to density, at least). The left panel shows the value of the statistic corresponding to the parameter BipartiteAltKCyclesB itself. The behaviour of this curve is starting to look as if it could be prone to near-degeneracy, with a fairly steep increase at a critical value.
The graphs in Fig. 1(b) show the results for the new BipartiteAltK4CyclesB parameter. In this case, both the Edge statistic (right) and BipartiteAltK4CyclesB statistic itself (left) show a phase transition, where a critical value of the parameter separates an empty graph regime from a complete graph regime. This model is therefore near-degenerate, suggesting that this new parameter may in fact be less useful than the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters.
For completeness, Fig. 1(c) shows the results for the new BipartiteFourCyclesNodePowerB parameter defined in the following section (Section 6). When $\alpha=0.5$ an abrupt change from a near-empty to a full graph occurs, however, as shown by the result for $\alpha=0.2$, this can be removed by decreasing the value of $\alpha$.
In summary, we defined new BipartiteAltK4CyclesA and BipartiteAltK4CyclesB parameters as modified forms of the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters defined in Wang et al. [21], in order to count four-cycles but not open two-paths. However, near-degeneracy was exhibited in the simulation experiments (similar results occur with the larger simulated networks described in Section 7; data not shown). We conclude, therefore, that although these new parameters could potentially be useful in some cases, they are more prone to near-degeneracy and not as useful as the existing $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters. Therefore, in the following section, we define new statistics that weight four-cycle counts in a different way.